OpenStudy (anonymous):

After swimming party at the Murrays house, there was only 200 2/3 gallons of water left in the pool (due to splashing). In order to refill the pool to its normal level, 25 7/18 gallons of water need to be added. What will be the full capacity of the pool after the new water is added? Reduce your answer to lowest terms.

OpenStudy (anonymous):

we started with: 200 2/3 gallons we added: 25 7/18 gallons when we add them together: 200 + 2/3 + 25 + 7/18 = 225 + 2/3 + 7/18 = 225 + (6*2)/(6*3) + 7/18 = 225 + 12/18 + 7/18 = 225 + 19/18 = 225 + 1 1/18 = 226 1/18

OpenStudy (anonymous):

we can do it a different way as well - but first let me know if the above is clear.

OpenStudy (anonymous):

yes that was right :) and clear thx <3

OpenStudy (anonymous):

above, I added the whole parts of the numbers seperately and the fractions seperately. You could also do it differently - by first converting to improper fractions, and adding those together.

OpenStudy (anonymous):

wow thanks for that :) that information help alot